# Vision Transformers for Multi-Variable Climate Downscaling: Emulating Regional Climate Models with a Shared Encoder and Multi-Decoder Architecture

### v1

@@ 2.2 Single-Variable Models
In the case of single variable the loss is calculated on the comparison between the generated output and the available ground truth, with our implementation using mean square error (MSE) loss. Denoting $\hat{Y}{hw}$ as the predicted pixel at spatial position $(h,w)$ and $Y{hw}$ as the corresponding ground truth pixel, the loss is defined as:
@@ 2.3 Multi-Variable Models
Both our proposed multi-variable approaches share the same loss structure. Training is performed simultaneously across the three target variables, with the loss computed as the mean over all output channels. To make this approach viable, it is essential to normalize all variables to the same range. This ensures that the model optimizes all variables with equal priority, preventing any one variable from dominating the loss due to scale differences. With MSE being our loss function of choice, and denoting $\hat{Y}{nhw}$ as the predicted pixel at spatial position $(h,w)$ for variable $n$, and $Y{nhw}$ as the corresponding ground truth pixel, the joint loss is defined as:
@@ 2.3.1 Single-Encoder Single-Decoder VIT (1E1D)
Our first approach to multi-variable modeling treats the different atmospheric fields as complementary features of a single spatial input by concatenating the variables along the channel dimension. This design keeps the single-variable architecture largely intact, with modifications only to the patch embedding and final convolutional layers to accommodate three input and output channels.
The input consists of three geophysical variables stacked as a three-channel image. We divide this input into non-overlapping patches of shape $8\times 8\times 3$, each of which is projected into a 256-dimensional vector via a dense embedding layer. These token embeddings are then processed by a Vision Transformer (ViT) encoder, following the same architecture used in the single-variable model. The decoder, composed of convolutional layers, upsamples the latent representation back to the original spatial resolution, with the final layer producing three output channels corresponding to the predicted fields.
This approach allows the encoder to capture a joint representation of the atmospheric state across multiple variables. However, it implicitly assumes that the variables are spatially aligned and semantically compatible when treated as separate channels, an assumption that may be limiting if the variables exhibit different spatial structures or temporal dynamics.
@@ 2.3.2 Single-Encoder Multi-Decoder VIT (1EMD)
The three target variables in our downscaling task exhibit substantial differences in their spatial structures and statistical distributions. For example, temperature fields tend to show smoother gradients compared to the more intricate patterns found in wind components. When using a single decoder to reconstruct all variables jointly, these differences can cause interference: the decoder is limited to sharing the same weights for generating the different variables, leading to potential cross-variable contamination. In particular, we observed that high-frequency patterns from one variable (e.g., wind) can bleed into the smoother reconstructions of another (e.g., geopotential height), degrading performance.
@@
Specifically, the 1EMD architecture is structured as follows. Each low-resolution input variable is first bilinearly upsampled to the target resolution of $504\times 432$, matching the output grid. The resulting variables are concatenated along the channel dimension and partitioned into 3402 non-overlapping patches of size $8\times 8\times 3$. Each patch is linearly embedded into a 256-dimensional vector and augmented with positional information via additive encoding. The sequence of embedded patches is processed by a stack of six transformer blocks, each featuring multi-head self-attention (six heads), residual connections with group normalization, and MLP layers with GELU activations.
The final encoder output is reshaped into a spatial feature grid of size $54\times 63\times 256$ and passed in parallel to three independent convolutional decoders, one per variable. The decoders do not share parameters and are trained jointly using a composite loss function that accounts for all variables. Each decoder consists of a sequence of residual blocks with group normalization and Swish activations, interleaved with 2D upsampling layers. Three upsampling stages are used to restore the original resolution of $504\times 432$. The final output is produced by a zero-initialized convolutional layer without activation. The complete 1EMD model architecture is reported in Figure 3.
@@ 3 Experimental Setup
We focus on downscaling three key atmospheric variables: near-surface air temperature (tas), surface wind speed (sfcWind), and 500 hPa geopotential height (zg500). From an image processing perspective, each variable is treated as a 2D field with spatial dimensions of $431\times 501$ pixels.
@@
Our downscaling models operate in an image-to-image manner, taking the GCM variable(s) at a given timestamp as input and predicting the corresponding RCM variable(s). In the single-variable case, the model maps one GCM field to one RCM field. In the multi-variable case, the model maps all three GCM variables to their corresponding RCM fields for the same timestamp, effectively forming a three-to-three mapping as illustrated in Figure 2.
We adopt a fair comparison framework that is coherent with standard practice in model comparison, avoiding common pitfalls such as unequal fine-tuning, training effort or model complexity [51, 52, 53]. We compare three modeling approaches, all based on a Vision Transformer (ViT) backbone. First, we train individual single-variable models for each of the three variables. Then, we explore two multi-variable models: 1E1D, which uses a shared encoder and decoder, and 1EMD, which employs a shared encoder and separate decoders for each variable.
All models share the same encoder architecture. The single-variable ViT and the multi-variable ViT 1E1D also share the same parameter count, differing only in the final convolutional layer and decoder embeddings. In contrast, the 1EMD model maintains a shared encoder but uses three independent decoders (one per variable), with no weight sharing in the decoding stage.
We use the Adam optimizer with decoupled weight decay (AdamW) [54], a common choice for ViT-based architectures. All models are trained for 400 epochs with 500 steps per epoch. Due to memory constraints, we use a batch size of 1. The learning rate is fixed at $1\mathrm{e}{-4}$ for the first 300 epochs and reduced to $1\mathrm{e}{-5}$ for the final 100 epochs. Weight decay is set to one-tenth of the current learning rate. These settings were chosen based on empirical tuning to ensure consistent convergence across all approaches. We use mean squared error (MSE) as the loss function for all models, as described in Sections 2.2 and 2.3.
@@ 4 Results
Figure 4: Comparison of the examined models (Single var ViT, multi var ViT 1E1D, Multi var ViT 1EMD), target RCM, and bilinearly interpolated GCM for two different samples of the test set. For each sample, each row contains geopotential, wind speed and temperature for the same timestamp.
GCM to RCM Downscaling Performance (MPI to RegCM)
@@
Bilinear 3.31e-03 4.82e-02 9.69e-01
Single Var. ViT 4.37e-04 1.47e-02 9.88e-01
Multi Var. ViT (1E1D) 4.63e-04 1.52e-02 9.84e-01
Multi Var. ViT (1EMD) 4.19e-04 1.43e-02 9.88e-01
@@
Bilinear 7.22e-03 6.40e-02 8.05e-01
Single Var. ViT 3.45e-03 4.37e-02 8.82e-01
Multi Var. ViT (1E1D) 3.31e-03 4.26e-02 8.79e-01
Multi Var. ViT (1EMD) 3.24e-03 4.21e-02 8.87e-01
@@
Single Var. ViT 1.89e-03 3.17e-02 9.95e-01
Multi Var. ViT (1E1D) 1.88e-03 3.15e-02 9.94e-01
Multi Var. ViT (1EMD) 1.81e-03 3.10e-02 9.95e-01
Table 1: Model performance comparison for downscaling from MPI-ESM-LR to RegCM4-6 for each variable. Best values for each variable are underlined, the results are computed for variables normalized between 0 and 1. Best results (including ties) are underlined.
@@
We evaluate the three models, Single-variable ViT, Multi-variable ViT (1E1D), and Multi-variable ViT (1EMD) on a test dataset comprising ten years of unseen data. The test set includes one year per decade from 2010 to 2100.
@@
All models are tested across the three selected variables: surface temperature (tas), wind speed (sfcWind), and 500 hPa geopotential height (zg500), using three evaluation metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Structural Similarity Index (SSIM). These metrics offer a well-rounded perspective, balancing numerical accuracy (MSE, MAE) with perceptual similarity (SSIM), and are commonly adopted in climate imaging studies [3, 55, 56, 57, 5].
All metrics are computed on normalized data to enable cross-variable comparison as the variables differ in their units and ranges.
Figure 5: Visualization of cross-variable leakage in the 1E1D model for zg500. From left to right: original GCM input, target RCM output, prediction from the 1E1D model, and prediction from the 1EMD model. While zg500 is expected to be spatially smooth, the 1E1D prediction exhibits faint outlines resembling European geography, suggesting contamination from temperature and wind fields. This artifact is absent in the 1EMD model, which treats each variable with dedicated decoder pathways.
Figure 6: Mean absolute error maps comparing the performance of bilinear interpolation (top row) and the multi-variable, multi-decoder 1EMD ViT model (bottom row). Each column shows the spatial distribution of errors for one variable: geopotential height at 500 hPa (zg500), 10-meter wind speed, and near-surface temperature. The error are computed for the complete test set, color bars are individually scaled per variable
Table 1 reports the numerical results for all models and variables. We also include bilinear interpolation as a baseline, applied directly to the GCM data before comparison with RCM targets. Figure 4 provides a visual comparison of all models alongside the original RCM data for two randomly selected test samples.
The Single-variable ViT consistently improves upon the bilinear interpolation baseline across all three variables. In the case of surface temperature (tas), the Single-variable ViT reduces MSE from 3.31e-03 to 4.37e-04 , and improves SSIM from 0.969 to 0.988. Similar improvements are observed for wind speed, where the MSE drops from 7.22e-03 to 3.45e-03, and for zg500, where it decreases from 2.55e-03 to 1.89e-03.
The Multi-variable ViT (1E1D) model improves over the Single-variable ViT for wind speed but performs worse on temperature and geopotential. This suggests that a single decoder doesn’t have sufficient capacity to model the three distinct variables simultaneously. The observed performance degradation indicates that the shared decoder architecture is likely a limiting factor, as it cannot fully exploit cross-variable information without compromising variable-specific accuracy. Additionally, the 1E1D model exhibits cross-variable leakage in zg500, with geographic artifacts emerging in the output, as illustrated in Figure 5.
@@
The Multi-variable ViT (1EMD) achieves the best results across all variables. Compared to the Single-variable ViT, For tas, MSE decreases from 4.37e-04 to 4.19e-04 while SSIM remains at 0.988. For wind speed, the MSE is further reduced from 3.45e-03 to 3.24e-03 and SSIM increases from 0.882 to 0.887. For zg500, MSE drops from 1.89e-03 to 1.81e-03. These results indicate that, under fair training conditions, a multi-task design with dedicated decoders is capable of leveraging cross-variable relationships to improve overall performance. To further assess the performance of 1EMD, a visual comparison of mean absolute error maps between 1EMD and bilinear interpolation is reported in Figure 6.
@@
Figure 7 presents the total runtime per model (in seconds), along with per-variable time for multi-variable models which process all three variables simultaneously.
The results show that, on a per-variable basis, the multi-variable models achieve approximately 25% faster inference compared to equivalent single-variable models.
Figure 7: Inference time comparison across model architectures. The Total inference time (dark blue bars) and per-variable inference time (light blue bars) are shown for three model configurations. All models were evaluated on identical hardware (NVIDIA A100), with batch size 1 and the same input data pipeline. The measurements correspond to the inference time required to process 10 years of daily testing data (3650 samples). For the multi-variable models, which predict three variables simultaneously, the per-variable inference time was calculated by dividing the total inference time by three.

### v2

@@ 2.2 Single-Variable Models
In the case of single variable the loss is calculated on the comparison between the generated output and the available ground truth, with our implementation using mean square error (MSE) loss. Denoting $\hat{Y}{hw}$ as the predicted pixel at spatial position $(h,w)$ and YhwY{hw} as the corresponding ground truth pixel, the loss is defined as:
@@ 2.3 Multi-Variable Models
Both our proposed multi-variable approaches share the same loss structure. Training is performed simultaneously across the six target variables, with the loss computed as the mean over all output channels. To make this approach viable, it is essential to normalize all variables to the same range. This ensures that the model optimizes all variables with equal priority, preventing any one variable from dominating the loss due to scale differences. With MSE being our loss function of choice, and denoting $\hat{Y}{nhw}$ as the predicted pixel at spatial position $(h,w)$ for variable $n$, and YnhwY{nhw} as the corresponding ground truth pixel, the joint loss is defined as:
@@ 2.3.1 Single-Encoder Single-Decoder VIT (1E1D)
Our first approach to multi-variable modeling treats the different atmospheric fields as complementary features of a single spatial input by concatenating the variables along the channel dimension. This design keeps the single-variable architecture largely intact, with modifications only to the patch embedding and final convolutional layers to accommodate six input and output channels.
The input consists of six atmospheric variables stacked as a six-channel image. We divide this input into non-overlapping patches of shape $8\times 8\times 6$, each of which is projected into a 256-dimensional vector via a dense embedding layer. These token embeddings are then processed by a Vision Transformer (ViT) encoder, following the same architecture used in the single-variable model. The decoder, composed of convolutional layers, upsamples the latent representation back to the original spatial resolution, with the final layer producing six output channels corresponding to the predicted fields.
This approach allows the encoder to capture a joint representation of the atmospheric state across multiple variables. However, it implicitly assumes that the variables are spatially aligned and compatible when treated as separate channels, an assumption that may be limiting if the variables exhibit different spatial structures or temporal dynamics.
@@ 2.3.2 Single-Encoder Multi-Decoder VIT (1EMD)
The six target variables in our downscaling task exhibit substantial differences in their spatial structures and statistical distributions. For example, temperature fields tend to show smoother gradients compared to the more intricate patterns found in wind components. When using a single decoder to reconstruct all variables jointly, these differences can cause interference: the decoder is limited to sharing the same weights for generating the different variables, leading to potential cross-variable contamination. In particular, we observed that high-frequency patterns from one variable (e.g., wind) can bleed into the smoother reconstructions of another (e.g., geopotential height, tp), degrading performance.
@@
Specifically, the 1EMD architecture is structured as follows. Each low-resolution input variable is first bilinearly upsampled to the target resolution of $504\times 432$, matching the output grid. The resulting variables are concatenated along the channel dimension and partitioned into 3402 non-overlapping patches of size $8\times 8\times 6$. Each patch is linearly embedded into a 256-dimensional vector and augmented with positional information via additive encoding. The sequence of embedded patches is processed by a stack of six transformer blocks, each featuring multi-head self-attention (six heads), residual connections with group normalization, and MLP layers with GELU activations.
The final encoder output is reshaped into a spatial feature grid of size $54\times 63\times 256$ and passed in parallel to six independent convolutional decoders, one per variable. The decoders do not share parameters and are trained jointly using a composite loss function that accounts for all variables. Each decoder consists of a sequence of residual blocks with group normalization and Swish activations, interleaved with 2D upsampling layers. Three upsampling stages are used to restore the original resolution of $504\times 432$. The final output is produced by a zero-initialized convolutional layer without activation. The complete 1EMD model architecture is reported in Figure 4.
@@
2.3.3 Multi-Variable U-net
In imaging tasks, the U-Net architecture represents a widely adopted standard baseline, with applications spanning much of the image-to-image translation spectrum. U-Nets have also been extensively used for statistical downscaling of meteorological variables, including several recent contributions specifically addressing multivariate downscaling [alipourhajiagha2025probabilistic, quesada2023downscaling, li2025generative]. To provide a meaningful and broadly recognized comparison against our transformer-based single-variable and multi-variable approaches, we include a multi-variable U-Net structured in a 1E1D configuration.
The implemented residual U-Net follows a standard symmetric encoder–decoder design built around residual blocks. Each residual block applies group normalization followed by two 3×3 convolutional layers. The encoder is composed of three downsampling stages, each containing three residual blocks and a 2× spatial reduction achieved via average pooling. Feature maps from every block are stored as skip connections and later concatenated in the decoder. The bottleneck stage processes the lowest-resolution representation through multiple residual blocks at the widest channel depth. The decoder mirrors the encoder, performing 2× bilinear upsampling at each stage and concatenating the corresponding skip features, followed by residual blocks
In the multi-variable, multi-task setting, the six variables are concatenated along the channel dimension in the input. A symmetric multi-channel output head is used to reconstruct the six variables, preserving a direct correspondence between input and output channels. The sizing of the model was designed to match the competing ViT based architectures in terms of execution and training times, The final configuration comprises approximately five million trainable parameters.
@@ 3 Experimental Setup
We focus on downscaling six key atmospheric variables: near-surface air temperature (tas), surface wind speed (sfcWind), 500 hPa geopotential height (zg500), total precipitation (tp), surface downwelling shortwave radiation (rsds), and surface downwelling longwave radiation (rlds). From an image processing perspective, each variable is treated as a 2D field with spatial dimensions of $431\times 501$ pixels.
@@
Our downscaling models operate in an image-to-image manner, taking the GCM variable(s) at a given timestamp as input and predicting the corresponding RCM variable(s). In the single-variable case, the model maps one GCM field to one RCM field. In the multi-variable case, the model maps all six GCM variables to their corresponding RCM fields for the same timestamp, effectively forming a six-to-six mapping as illustrated in Figure 3.
We adopt a fair comparison framework that is coherent with standard practice in model comparison, avoiding common pitfalls such as unequal fine-tuning, training effort or model complexity [lucic2018gans, pineau2021improving, lipton2019troubling]. We compare the four modeling approaches, three based on a Vision Transformer (ViT) backbone and one based on a U-Net. First, we train and test six independent single-variable ViT models, one for each target variable. Then, we explore three multi-variable models: 1E1D, which uses a shared encoder and decoder, 1EMD, which employs a shared encoder and separate decoders for each variable, and a multi-variable U-Net providing an external, widely used reference architecture. All ViT models share the same encoder architecture. The single-variable ViT and the multi-variable ViT 1E1D also share the same parameter count, differing only in the final convolutional layer and decoder embeddings. In contrast, the 1EMD model maintains a shared encoder but uses six independent decoders (one per variable), with no weight sharing in the decoding stage.
We use the Adam optimizer with decoupled weight decay (AdamW) [loshchilov2017decoupled], a common choice for ViT-based architectures. All models are trained for 400 epochs with 500 steps per epoch. Due to memory constraints, we use a batch size of 1. The learning rate is fixed at $1\mathrm{e}{-4}$ for the first 300 epochs and reduced to $1\mathrm{e}{-5}$ for the final 100 epochs. Weight decay is set to one-tenth of the current learning rate. These settings were chosen based on empirical tuning to ensure consistent convergence across all approaches. We use mean squared error (MSE) as the loss function for all models, as described in Sections 2.2 and 2.3.
@@
To account for variability due to random initialization and stochastic optimization, all models were trained and evaluated twice using independent random seeds. Each repetition executed the full training and testing pipeline under identical settings, and the performance values reported in the manuscript correspond to the mean across the two runs. This protocol provides a lightweight but meaningful assessment of robustness within realistic computational constraints. Notably, the relative ranking of the models is consistent across both repetitions.
@@ 4 Results
Figure 5: Comparison of the examined models (Single var ViT, multi var ViT 1E1D, Multi var ViT 1EMD), target RCM, and bilinearly interpolated GCM for a sample of the test set. For each sample, each row contains Geopotential, Wind speed, Temperature, Total Precipitation, Surface Downwelling Shortwave Radiation and Surface Downwelling Longwave Radiation for the same timestamp.
Mean Performance Across Two Independent Runs
@@
Bilinear 9.97e-04 2.27e-02 9.78e-01
Single Var. ViT 4.34e-04 1.47e-02 9.89e-01
Multi Var. ViT (1E1D) 4.50e-04 1.51e-02 9.84e-01
Multi Var. U-Net 4.47e-04 1.48e-02 9.83e-01
Multi Var. ViT (1EMD) 4.23e-04 1.43e-02 9.88e-01
@@
Bilinear 6.08e-03 5.81e-02 8.35e-01
Single Var. ViT 3.43e-03 4.36e-02 8.85e-01
Multi Var. ViT (1E1D) 3.33e-03 4.29e-02 8.76e-01
Multi Var. U-Net 3.35e-03 4.32e-02 8.69e-01
Multi Var. ViT (1EMD) 3.23e-03 4.21e-02 8.87e-01
@@
Single Var. ViT 1.90e-03 3.18e-02 9.95e-01
Multi Var. ViT (1E1D) 1.85e-03 3.14e-02 9.94e-01
Multi Var. U-Net 1.84e-03 3.17e-02 9.95e-01
Multi Var. ViT (1EMD) 1.80e-03 3.08e-02 9.95e-01
tp
Bilinear 6.00e-07 3.73e-04 9.97e-01
Single Var. ViT 4.46e-07 3.55e-04 9.97e-01
Multi Var. ViT (1E1D) 4.91e-07 3.94e-04 9.96e-01
Multi Var. U-Net 5.00e-07 4.09e-04 9.96e-01
Multi Var. ViT (1EMD) 4.32e-07 3.48e-04 9.97e-01
rsds
Bilinear 2.24e-02 1.01e-01 7.79e-01
Single Var. ViT 5.37e-03 5.02e-02 8.80e-01
Multi Var. ViT (1E1D) 5.03e-03 4.86e-02 8.76e-01
Multi Var. U-Net 5.00e-03 4.85e-02 8.75e-01
Multi Var. ViT (1EMD) 4.99e-03 4.83e-02 8.83e-01
rlds
Bilinear 6.70e-03 6.37e-02 9.17e-01
Single Var. ViT 3.06e-03 4.28e-02 9.43e-01
Multi Var. ViT (1E1D) 2.80e-03 4.03e-02 9.43e-01
Multi Var. U-Net 2.81e-03 4.05e-02 9.40e-01
Multi Var. ViT (1EMD) 2.78e-03 4.01e-02 9.47e-01
Table 1: Mean model performance across two independent runs. Best values (including ties) are underlined.
@@
We evaluate the four models, Single-variable ViT, Multi-variable ViT (1E1D), Multi-variable ViT (1EMD) and Multi-variable U-Net on a test dataset comprising ten years of unseen data. The test set includes one year per decade from 2010 to 2100.
@@
All models are tested across the six selected variables using three evaluation metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Structural Similarity Index (SSIM) alongside spatial and temporal Pearson correlations. These metrics offer a well-rounded perspective, balancing numerical accuracy (MSE, MAE) with perceptual similarity (SSIM), and are commonly adopted in climate imaging studies [bano2020configuration, asperti2025precipitation, yang2024fourier, harder2023hard, oyama2023deep]. All metrics are computed on normalized data to enable cross-variable comparison as the variables differ in their units and ranges.
Figure 6: Visualization of cross-variable leakage in the 1E1D and multi-variable U-Net models for total precipitation and zg500. Four samples are shown, two for precipitation and two for geopotential height. Using a high contrast colormap, we can clearly see high-frequency artefacts in the multi-variable U-Net and 1E1D ViT, with faint outlines resembling European geography. This suggests contamination from temperature and wind fields. These outlines are absent from the original data, the single-variable model, and the multi-decoder 1EMD model.
Spatial and Temporal Correlations (MPI to RegCM)
Model Spatial corr $(\mu\pm\sigma)$ Temporal corr $(\mu\pm\sigma)$
tas
Bilinear $0.910\pm 0.037$ $0.935\pm 0.016$
Single Var. ViT $0.957\pm 0.024$ $0.956\pm 0.011$
Multi Var. ViT (1E1D) $0.957\pm 0.024$ $0.956\pm 0.010$
Multi Var. U-Net $0.955\pm 0.024$ $0.956\pm 0.011$
Multi Var. ViT (1EMD) $\underline{0.958\pm 0.024}$ $\underline{0.957\pm 0.010}$
sfcWind
Bilinear $0.602\pm 0.161$ $0.528\pm 0.096$
Single Var. ViT $0.748\pm 0.113$ $0.605\pm 0.083$
Multi Var. ViT (1E1D) $0.759\pm 0.109$ $0.636\pm 0.075$
Multi Var. U-Net $0.755\pm 0.111$ $0.631\pm 0.081$
Multi Var. ViT (1EMD) $\underline{0.765\pm 0.109}$ $\underline{0.644\pm 0.073}$
zg500
Bilinear $0.929\pm 0.089$ $0.920\pm 0.033$
Single Var. ViT $0.943\pm 0.073$ $0.935\pm 0.024$
Multi Var. ViT (1E1D) $0.943\pm 0.070$ $0.937\pm 0.023$
Multi Var. U-Net $\underline{0.944\pm 0.070}$ $\underline{0.938\pm 0.023}$
Multi Var. ViT (1EMD) $\underline{0.944\pm 0.070}$ $0.937\pm 0.023$
tp
Bilinear $0.339\pm 0.173$ $0.340\pm 0.111$
Single Var. ViT $0.411\pm 0.158$ $0.389\pm 0.112$
Multi Var. ViT (1E1D) $0.341\pm 0.121$ $0.321\pm 0.116$
Multi Var. U-Net $0.293\pm 0.142$ $0.298\pm 0.128$
Multi Var. ViT (1EMD) $\underline{0.438\pm 0.150}$ $\underline{0.407\pm 0.106}$
rsds
Bilinear $0.745\pm 0.165$ $0.826\pm 0.048$
Single Var. ViT $0.885\pm 0.093$ $0.946\pm 0.016$
Multi Var. ViT (1E1D) $\underline{0.894\pm 0.088}$ $\underline{0.951\pm 0.014}$
Multi Var. U-Net $0.893\pm 0.088$ $0.950\pm 0.013$
Multi Var. ViT (1EMD) $\underline{0.894\pm 0.087}$ $\underline{0.951\pm 0.014}$
rlds
Bilinear $0.486\pm 0.168$ $0.703\pm 0.078$
Single Var. ViT $0.791\pm 0.089$ $0.811\pm 0.053$
Multi Var. ViT (1E1D) $\underline{0.808\pm 0.086}$ $\underline{0.831\pm 0.054}$
Multi Var. U-Net $0.805\pm 0.088$ $0.830\pm 0.055$
Multi Var. ViT (1EMD) $\underline{0.808\pm 0.088}$ $0.829\pm 0.054$
Table 2: Spatial and temporal Pearson correlations between model predictions and RegCM4-6 targets for each variable. Spatial correlations are computed for each day and summarized as mean $\pm$ standard deviation over time. Temporal correlations are computed at each grid point and summarized as mean $\pm$ standard deviation over space. Best values (including ties) are underlined.
Figure 7: Mean absolute error maps for bilinear interpolation (top rows) and the 1EMD multi-variable ViT (bottom rows). The top panel shows zg500, wind speed, and temperature; the bottom panel shows precipitation, shortwave downwelling radiation, and longwave downwelling radiation. Errors are computed over the full test set, with per-variable color scales.
Table 1 reports the mean performance across two independent runs for all six variables. All deep-learning models provide substantial improvements over the bilinear interpolation baseline. For example, the MSE for near-surface temperature (tas) decreases from $9.97\times 10^{-4}$ under bilinear interpolation to $4.34\times 10^{-4}$ with the Single-variable ViT. Similarly, for surface wind speed (sfcWind), the bilinear error of $6.08\times 10^{-3}$ is reduced by approximately half across all learned models. Comparable improvements are observed for the remaining variables, confirming the value of learning-based downscaling across a diverse set of atmospheric fields.
A second consistent pattern is that the Multi-var 1E1D ViT and the Multi-var U-Net achieve nearly identical results for all variables. For instance, on rsds the two models obtain MSE values of $5.03\times 10^{-3}$ and $5.00\times 10^{-3}$, while on rlds they reach $2.80\times 10^{-3}$ and $2.81\times 10^{-3}$, respectively. This close alignment suggests that architectural differences between transformer-based and convolutional backbones are less influential than the broader design choice of using a single shared decoder when performing multi-variable downscaling.
The comparison with the Single-variable ViT indicates that joint modeling through a single decoder results in positive transfer for several variables. Both 1E1D and the U-Net improve upon the Single-variable ViT for sfcWind (reducing MSE from $3.43\times 10^{-3}$ to $3.33\times 10^{-3}$ and $3.35\times 10^{-3}$), for zg500 (from $1.90\times 10^{-3}$ to $1.85\times 10^{-3}$ and $1.84\times 10^{-3}$), for rsds (from $5.37\times 10^{-3}$ to $5.03\times 10^{-3}$ and $5.00\times 10^{-3}$), and for rlds (from $3.06\times 10^{-3}$ to $2.80\times 10^{-3}$ and $2.81\times 10^{-3}$). These results indicate that several variables benefit from shared representation learning across related atmospheric patterns. In contrast, the Single-variable ViT performs better for tas and total precipitation (tp). For tas, the MSE increases from $4.34\times 10^{-4}$ when using single-variable modeling to $4.50\times 10^{-4}$ (1E1D) and $4.47\times 10^{-4}$ (U-Net). For tp, the error increases from $4.46\times 10^{-7}$ to $4.91\times 10^{-7}$ and $5.00\times 10^{-7}$. These results illustrate negative transfer for variables whose statistical or spatial characteristics are not well accommodated by a shared decoder. Additionally, the 1E1D model exhibits cross-variable leakage in zg500, with geographic artifacts emerging in the output, as illustrated in Figure 6.
@@
The proposed 1EMD architecture achieves the best performance for all six variables. Relative to the Single-variable ViT, the MSE decreases from $4.34\times 10^{-4}$ to $4.23\times 10^{-4}$ for tas, from $3.43\times 10^{-3}$ to $3.23\times 10^{-3}$ for sfcWind, and from $4.46\times 10^{-7}$ to $4.32\times 10^{-7}$ for tp. The improvements are consistent across the remaining variables as well. These results indicate that combining a shared encoder with variable-specific decoders successfully incorporates the benefits of multi-variable modeling while mitigating the negative transfer observed in single-decoder approaches. To further assess the performance of 1EMD, a visual comparison of mean absolute error maps between 1EMD and bilinear interpolation is reported in Figure 7.
Additionally, Table 2 reports spatial and temporal Pearson correlations for all variables and models. Spatial correlation quantifies, at each time step, the agreement between predicted and observed spatial patterns and is summarized as mean ± standard deviation across time. Temporal correlation analogously measures, at each grid point, the agreement between predicted and observed time series and is summarized across space. Across variables, the spatial and temporal correlations reveal two broad regimes of model behaviour. For tas and zg500, all learned models perform similarly, with minimal differences between the single-variable and multivariate configurations. For rsds and rlds, learned models also converge to closely aligned performance, though a modest improvement is still observed when moving from single-variable to multivariate setups. In contrast, wind speed and total precipitation (tp) exhibit much larger performance spreads. For wind speed, the advantage of 1EMD is substantial: it clearly outperforms both the single-variable model and the other multivariate variants, achieving the highest spatial and temporal correlations (up to 0.76 and 0.64, respectively), whereas the alternative multivariate designs repeatedly underperform. Precipitation displays even more pronounced differences: most multivariate models deteriorate markedly (e.g., temporal correlation dropping to 0.32 in 1E1D), while 1EMD reaches 0.41, surpassing both the single-variable model (0.39) and all other multivariate approaches. Overall, 1EMD emerges as the most robust multivariate design, with the largest gains observed for the more challenging variables.
@@
In Table 3 we report the number of parameters, execution time, and training time for all models. Although the 1EMD model has a substantially larger number of trainable parameters due to its six independent decoders, its inference time remains similar to that of the 1E1D model. This is expected: using six decoders instead of one increases the parameter count but does not significantly change the number of arithmetic operations, as each decoder performs the same sequence of computations with different weights. Moreover, the transformer encoder dominates the total computational cost, meaning that differences in decoder size contribute only marginally to overall runtime.
Figure 8 presents the total runtime per model (in seconds), along with per-variable time for multi-variable models which process all six variables simultaneously.
[…]
Figure 8: Inference time comparison across model architectures. The Total inference time (dark blue bars) and per-variable inference time (light blue bars) are shown for four model configurations. All models were evaluated on identical hardware (NVIDIA A100), with the same input data pipeline. The measurements correspond to the inference time required to process 10 years of daily testing data. For the multi-variable models, which predict six variables simultaneously, the per-variable inference time was calculated by dividing the total inference time the number of variables.
Model Exec. Time (s) Train Time (s) Params (M)
Single-variable ViT 817 11 200 11.63
1E1D ViT 3 409 58 400 15.39
Multi-variable U-Net 3 336 61 320 5.20
1EMD ViT 3 456 60 200 32.25
Table 3: Computational cost and model complexity for the three examined ViT architectures. Execution time is measured over the entire 10-year test dataset. Training time corresponds to 400 training epochs under identical hardware and hyperparameter settings.